solve the following system by substitution.


4x -12y= 5
-x +3y=-1

Answer by venugopalramana(3286)

4x -12y= 5..............I
-x +3y=-1..............II
FROM EQN.II
X=1+3Y..PUT IN EQN.I
4(1+3Y)-12Y=5
4+12Y-12Y=5
4=5..NOT POSSIBLE.
THERE IS NO SOLUTION AS THE EQNS.ARE INCONSISTENT